A body is projected horizontally from a point above the ground. the motion of the body  is described  the equations x=2  and y=5t^2, where x and y are the horizontal and vertical displacements (in m) respectively at time t. What is the magnitude of the velocity of the body 0.2 seconds after it is projected?​

Grade:11

1 Answers

Aditya Gupta
4 years ago
we see that x coordinate remains constant. which means that the flight of the object is vertical. so, its velocity will only have a vertical component.
now, v= dy/dt= 10t
at t= 0.2, v= 10*0.2= 2m/s
kindly approve :)